OpenStudy (anonymous):

Solve x+2/15 + 3-x/5=2/3

OpenStudy (lgbasallote):

\[\huge \frac{x+2}{15} + \frac{3-x}{5} = \frac{2}{3}\] hint: use 15 as LCM. multiply each term by 15

OpenStudy (anonymous):

@lgbasallote idk how to do this >:(

OpenStudy (lgbasallote):

what do you mean?

OpenStudy (anonymous):

i dont know how to solve this at all.

OpenStudy (lgbasallote):

but i just told you how..i don't get what you mean...multiply 15 to each term..have you tried doing it?

OpenStudy (anonymous):

wouldnt i multiply each term by either 5 or 3 to get a common denominator as 3?

OpenStudy (anonymous):

common denominator of 15 my bad